Find the minimum value of g(x)=5x^2+17x-3
Answer by nerdybill(7384)

g(x)=5x^2+17x-3
.
Because it is in the form of
Ax^2 + Bx + C
it is a "quadratic" -- a parabola
By looking at the coefficient associated with the x^2 term we can tell whether it opens upwards (positive coefficient) or downwards (negative coefficient). It is positive so it opens upwards.
.
Finding the vertex give you the minimum.
it is minimum when
x = -b/(2a) = -17/(2*5) = -17/10 = -1.7
.
Plug it back into the original equation to find the minimum:
g(x)=5x^2+17x-3
g(-1.7)=5(-1.7)^2+17(-1.7)-3
g(-1.7)= -17.45